A rectangular storage area is 8.0 m longer than it is wide. If the area is 28 m squared, what are its dimensions?
I have such trouble with word problems!

Answer by ahenkemeyer(12)

Area = Length * Width
A = L*W
.
We're told
A = 28
.
We're also told the length is 8 meters greater than width.
L = W+8
.
L * W = A
.
Substituting:
(W+8) * W = 28
W^2 + 8W = 28
Subtract 28 from both sides
W^2 + 8W - 28 = 0